In the experiment for the moment of inertia of the
Answer #1


alpha = a / R


alpha = 0.5 / (15 x 10^-3)

alpha = 33.33 rad/s^2 ...... Angular acc of disk.


Applying net torque = I alpha

{ torque = r T}

(15 x 10^-3) (5) = I (33.33)

I = 2.25 x 10^-3 kg m^2